$SPAD/src/input d02bhf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
xend:SF:=10.0
n:=3
irelab:=0
hmax:SF:=0.0
x:SF:=0.0
y:Matrix SF:=
 [[0.5 ,0.5 ,%pi*0.2 ]]
tol:SF:=0.0001
ef:Expression Float:=1.0*Y[1]::EXPR FLOAT
g:Asp9(G):=retract ef
vef:Vector Expression Float:=
 [tan(Y[3]) ,-0.032*tan(Y[3])/Y[2] -0.02*Y[2]/cos(Y[3]) ,-0.032/(Y[2]**2) ]
fcn:Asp7(FCN):= retract vef
result:=d02bhf(xend,n,irelab,hmax,x,y,tol,-1,g,fcn)

———-

References^{[1] nothing}

4